# Test-time Adaptation for Regression by Subspace Alignment

### v1

@@ 1 Introduction
Deep neural networks have achieved remarkable success in various tasks [35, 33, 23, 10]. In particular, regression, which is one of the fundamental tasks in machine learning, is widely used in practical tasks such as human pose estimation or age prediction [34]. The successes of deep learning have usually relied on the assumption that the training and test datasets are sampled from an i.i.d. distribution. In the real world, however, such an assumption is often invalid since the test data are sampled from distributions different from the training one due to distribution shifts caused by changes in environments. The performance of these models thus deteriorates when a distribution shift occurs [25, 50]. To address this problem, test-time adaptation (TTA) [40] has been studied. TTA aims at adapting a model pre-trained in a source domain (training environment) to the target domain (test environment) with only unlabeled target data. However, most of the existing TTA methods are designed for classification; that is, TTA for regression has not been explored much [40]. Regarding TTA for classification, two main approaches have been explored: entropy minimization and feature alignment.
@@
The entropy minimization approach was introduced by Wang et al. [54], and the subsequent methods follow this approach [61, 46, 57, 60, 17]. Although entropy is a promising proxy of the performance on the target domain, entropy minimization is classification-specific because it assumes that a model directly outputs predictive distributions, i.e., a probability for each class. In contrast, typical regression models output only single scalar values, not distributions. Thus, we cannot use the entropy minimization approach for regression models.
@@
Another approach, the feature alignment, preliminarily computes the statistics of intermediate features of the source dataset after pre-training in the source domain [27, 31, 16, 1, 29]. Then, upon moving to the target domain, the feature distribution of the target data is aligned with the source distribution by matching the target feature statistics with the pre-computed source ones without accessing the source dataset. Although this approach seems applicable to regression because it allows arbitrary forms of the model output, it assumes to use all feature dimensions to be aligned, and does not sufficiently consider the nature of regression tasks. For instance, regression models trained with standard mean squared error (MSE) loss tend to make features less diverse than classification models do [58]. In particular, we experimentally observed that the features of a trained regression model are distributed in only a small subspace of the entire feature space (Tab. 1). In this sense, naively aligning all feature dimensions makes the performance suboptimal or even be harmful in regression as shown in Tab. 1 because it equally treats important feature dimensions and degenerated unused ones.
@@
We conducted experiments on various regression tasks, such as UTKFace [59], Biwi Kinect [18], and California Housing [47]. The results showed that our SSA retains the important feature subspace during TTA and outperforms existing TTA baselines that were originally designed for classification by aligning the feature subspace.
Table 1: Number of valid (having non-zero variance) feature dimensions and feature subspace dimensions (i.e., the rank of the feature covariance matrix), and $R^{2}$ scores on the test datasets. Although the original feature space has 2048 dimensions in the experiment except for the California Housing dataset, the features of the regression models are distributed within the subspaces that have less than a hundred dimensions. Our method improves $R^{2}$ scores by the feature subspace in most cases, whereas the naive feature alignment sometimes diverged (displayed as ‘-’) because of too few valid dimensions. See Sec. 4.3.1 for more details.
@@
Dataset #Valid dims #Subspace dims. Source Naive feature alignment SSA (ours)
@@
Biwi Kinect (mean) 713 34.5 $0.706$ $0.753$ $\mathbf{0.778}$
@@ 2 Problem Setting
We consider a setting with a neural network regression model $f_{\theta}:\mathcal{X}\to\mathbb{R}$ pre-trained on a labeled source dataset $\mathcal{S}=\{(\mathbf{x}_{i}^{\text{s}},y_{i}^{\text{s}})\in\mathcal{X}\times\mathbb{R}\}_{i=1}^{N_{\text{s}}}$, where $\mathbf{x}_{i}^{\text{s}}$ and $y_{i}^{\text{s}}$ are an input and its label, and $\mathcal{X}$ is the input space. Our goal is to adapt $f_{\theta}$ to the target domain by using an unlabeled target dataset $\mathcal{T}=\{\mathbf{x}_{i}^{\text{t}}\in\mathcal{X}\}_{i=1}^{N_{\text{t}}}$ without accessing $\mathcal{S}$. Note that the target labels $y_{i}^{\text{t}}\in\mathbb{R}$ are not available. In the source dataset $\mathcal{S}$, the data $\{(\mathbf{x}_{i}^{\text{s}},y_{i}^{\text{s}})\}$ are sampled from the source distribution $p_{\text{s}}$ over $\mathcal{X}\times\mathbb{R}$. In the target dataset $\mathcal{T}$, we assume covariate shift [52], which is a distribution shift that often occurs in the real world. In other words, the target data $\mathbf{x}_{i}^{\text{t}}$ are sampled from the target distribution $p_{\text{t}}$ over $\mathcal{X}$ that is different from $p_{\text{s}}$, but the predictive distribution is the same, i.e., $p_{\text{s}}(\mathbf{x})\neq p_{\text{t}}(\mathbf{x})$ and $p_{\text{s}}(y|\mathbf{x})=p_{\text{t}}(y|\mathbf{x})$.
@@ 3 Test-time Adaptation for Regression
Figure 1: Overview of significant-subspace alignment (SSA).
In this section, we describe the basic idea behind Significant-subspace Alignment (SSA) in Sec. 3.1 and describe it in detail in Sec. 3.2. Our method can be applied regardless of the form of input data since SSA does not rely on input-specific method, such as image data augmentations or self-supervised tasks.
@@ 3.1 Basic Idea: Feature Alignment
$\bm{\mu}^{\text{s}}$ $=\frac{1}{N_{\text{s}}}\sum_{i=1}^{N_{\text{s}}}\mathbf{z}_{i}^{\text{s}},$ (1)
$\bm{\sigma}^{\text{s~{}2}}$ $=\frac{1}{N_{\text{s}}}\sum_{i=1}^{N_{\text{s}}}(\mathbf{z}_{i}^{\text{s}}-\bm{\mu}^{\text{s}})\odot(\mathbf{z}_{i}^{\text{s}}-\bm{\mu}^{\text{s}}),$ (2)
@@
Then, we move to the target domain, where we cannot access the source dataset $\mathcal{S}$. Given a target mini-batch =$\mathcal{B}=\{\mathbf{x}_{i}^{\text{t}}\}_{i=1}^{B}$ sampled from $\mathcal{T}$, we compute the mini-batch mean and variance $\hat{\bm{\mu}}^{\text{t}}$ and $\hat{\bm{\sigma}}^{\text{t~{}2}}$ analogously to Eqs. 1 and 2.
For feature alignment, we seek to make the target statistics similar to the source ones. For this purpose, we use the KL divergence as Nguyen et al. [45] proved that it is included in an upper bound of the target error in unsupervised domain adaptation. Concretely, we minimize the KL divergence between two diagonal Gaussian distributions $\mathcal{N}(\bm{\mu}^{\text{s}},{\bm{\sigma}^{\text{s~{}2}}})$ and $\mathcal{N}(\hat{\bm{\mu}}^{\text{t}},\hat{\bm{\sigma}}^{\text{t~{}2}})$:
$\mathcal{L}_{\text{TTA}}(\phi)$ $=\sum_{d=1}^{D}D_{\text{KL}}\left(\mathcal{N}(\mu_{d}^{\text{s}},{{\sigma}_{d}^{\text{s}}}^{2})\|\mathcal{N}(\hat{\mu}_{d}^{\text{t}},{\hat{{\sigma}}_{d}^{\text{t~{}2}}})\right)$
$\hskip 17.50002pt+D_{\text{KL}}\left(\mathcal{N}(\hat{\mu}_{d}^{\text{t}},{\hat{{\sigma}}_{d}^{\text{t~{}2}}})\|\mathcal{N}(\mu_{d}^{\text{s}},{{\sigma}_{d}^{\text{s}}}^{2})\right),$ (3)
where the subscripts $d$ represent the $d$-th elements of the mean and variance vectors. Here, we used both directions of the KL divergence because it empirically had good results as recommended by Nguyen et al. [45]. The KL divergence between two univariate Gaussians can be written in a closed form [15]
$D_{\text{KL}}$ $\left(\mathcal{N}(\mu_{1},\sigma_{1}^{2})\|\mathcal{N}(\mu_{2},\sigma_{2}^{2})\right)$
$\hskip 30.00005pt=\frac{1}{2}\left[\log\frac{\sigma_{2}^{2}}{\sigma_{1}^{2}}+\frac{(\mu_{1}-\mu_{2})^{2}+\sigma_{1}^{2}}{\sigma_{2}^{2}}-1\right].$ (4)
However, in regression models, the features tend to be less diverse than in classification [58]. In addition to the insight from Zhang et al. [58], we observed that the features of a regression model trained on $\mathcal{S}$ are distributed in only a small subspace of the feature space and many dimensions of the feature space had zero variances. Tab. 1 shows the numbers of valid (having non-zero variance) feature dimensions and feature subspace dimensions (see Sec. 4.3.1 for more details). This property makes the naive feature alignment described above unstable since the KL divergence in Eq. 4 includes the variance in the denominator. Also, this naive feature alignment is ineffective because many of the feature dimensions have a small effect on the subspace.
@@ 3.2 Significant-subspace Alignment
In this section, we describe our method, Significant-subspace Alignment (SSA), to tackle the aforementioned problem of naive feature alignment. Fig. 1 shows an overview of SSA. As described in Sec. 3.1, the features of a regression model tend to be distributed in a small subspace of the feature space. Thus, we introduce subspace detection to detect a subspace that is representative and significant to the output and then perform feature alignment in the subspace. Subspace detection is similar to principal component analysis (PCA). Further, in our regression model, a feature $\mathbf{z}=g_{\phi}(\mathbf{x})$ is projected onto a one-dimensional line determined by $\mathbf{w}$ of $h_{\psi}$ in order to output a scalar value $\mathbf{w}^{\top}\mathbf{z}+b$. Thus, a subspace basis $\mathbf{v}^{\text{s}}_{d}$ whose direction is not orthogonal to $\mathbf{w}$ needs precise feature alignment. We use dimension weighting to prioritize such dimensions.
Subspace detection. After the training on the source dataset $\mathcal{S}$, we detect the subspace in which the source features are distributed. Instead of computing the variance of each dimension as Eqs. 1 and 2, we compute the covariance matrix:
@@
where the source mean vector $\bm{\mu}^{\text{s}}$ is the same as Eqs. 1 and 2.
@@
where $\mathbf{V}^{\text{s}}=[\mathbf{v}_{1}^{\text{s}},\ldots,\mathbf{v}_{K}^{\text{s}}]^{\top}\in\mathbb{R}^{K\times D}$. With $\tilde{\mathbf{z}}^{\text{t}}_{i}\in\mathbb{R}^{K}$, we compute the projected target mean and variance over the mini-batch analogously to Eqs. 1 and 2; this is denoted by ${\tilde{\bm{\mu}}^{\text{t}}}$ and $\tilde{\bm{\sigma}}^{\text{t~{}2}}$. On the other hand, the projected source mean and variance are $\mathbf{0}$ and the eigenvalues $\bm{\lambda}^{\text{s}}=[\lambda_{1}^{\text{s}},\ldots,\lambda_{K}^{\text{s}}]$ since $\bm{\Sigma^{\text{s}}}\mathbf{v}_{k}^{\text{s}}=\lambda^{\text{s}}_{k}\mathbf{v}_{k}^{\text{s}}$. Thus, the KL divergence in the detected subspace is computed between two $K$-dimensional diagonal Gaussians $\mathcal{N}(\mathbf{0},\bm{\lambda}^{\text{s}})$ and $\mathcal{N}({\tilde{\bm{\mu}}^{\text{t}}},\tilde{\bm{\sigma}}^{\text{t~{}2}})$. Using subspace detection and dimension weighting, the loss of SSA is:
$\mathcal{L}_{\text{TTA}}(\phi)$ $=\sum_{d=1}^{K}\alpha_{d}\left\{D_{\text{KL}}\left(\mathcal{N}(0,\lambda_{d}^{\text{s}})\|\mathcal{N}({\tilde{\mu}_{d}^{\text{t}}},{\tilde{\sigma}_{d}^{\text{t~{}2}}})\right)\right.$
$\hskip 37.00002pt\left.+D_{\text{KL}}\left(\mathcal{N}({\tilde{\mu}_{d}^{\text{t}}},{\tilde{\sigma}_{d}^{\text{t~{}2}}})\|\mathcal{N}(0,\lambda_{d}^{\text{s}})\right)\right\}$
@@
During TTA, we optimize the feature extractor $g_{\phi}$ to minimize $\mathcal{L}_{\text{TTA}}$, i.e., we seek $\phi^{*}=\min_{\phi}\mathcal{L}_{\text{TTA}}(\phi)$. We update only the affine parameters $\gamma$ and $\beta$ of the normalization layers such as batch normalization [26] or layer normalization [2] inspired by Tent [54]. This strategy is effective not only to avoid forgetting the source knowledge but also to enable flexible adaptation [20, 5]. The procedure of SSA is listed in Algorithm 1 of the Appendix.

### v2

@@ 1 Introduction
Deep neural networks have achieved remarkable success in various tasks (LeCun et al., 1998a; Krizhevsky et al., 2012; He et al., 2016; Dargan et al., 2020). In particular, regression, which is one of the fundamental tasks in machine learning, is widely used in practical tasks such as human pose estimation or age prediction (Lathuilière et al., 2019). The successes of deep learning have usually relied on the assumption that the training and test datasets are sampled from an i.i.d. distribution. In the real world, however, such an assumption is often invalid since the test data are sampled from distributions different from the training one due to distribution shifts caused by changes in environments. The performance of these models thus deteriorates when a distribution shift occurs (Hendrycks & Dietterich, 2019; Recht et al., 2019). To address this problem, test-time adaptation (TTA) (Liang et al., 2023) has been studied. TTA aims at adapting a model pre-trained in a source domain (training environment) to the target domain (test environment) with only unlabeled target data. However, most of the existing TTA methods are designed for classification; that is, TTA for regression has not been explored much (Liang et al., 2023). Regarding TTA for classification, two main approaches have been explored: entropy minimization and feature alignment.
@@
The entropy minimization approach was introduced by Wang et al. (2021), and the subsequent methods follow this approach (Zhou & Levine, 2021; Niu et al., 2022; Zhang et al., 2022; Zhao et al., 2023; Enomoto et al., 2024). Although entropy is a promising proxy of the performance on the target domain, entropy minimization is classification-specific because it assumes that a model directly outputs predictive distributions, i.e., a probability for each class. In contrast, typical regression models output only single scalar values, not distributions. Thus, we cannot use the entropy minimization approach for regression models.
@@
Another approach, the feature alignment, preliminarily computes the statistics of intermediate features of the source dataset after pre-training in the source domain (Ishii & Sugiyama, 2021; Kojima et al., 2022; Eastwood et al., 2022; Adachi et al., 2023; Jung et al., 2023). Then, upon moving to the target domain, the feature distribution of the target data is aligned with the source distribution by matching the target feature statistics with the pre-computed source ones without accessing the source dataset. Although this approach seems applicable to regression because it allows arbitrary forms of the model output, it assumes to use all feature dimensions to be aligned, and does not sufficiently consider the nature of regression tasks. For instance, regression models trained with standard mean squared error (MSE) loss tend to make features less diverse than classification models do (Zhang et al., 2023). In particular, we experimentally observed that the features of a trained regression model are distributed in only a small subspace of the entire feature space (Table 1). In this sense, naively aligning all feature dimensions makes the performance suboptimal or even be harmful in regression as shown in Table 1 because it equally treats important feature dimensions and degenerated unused ones.
@@
We conducted experiments on various regression tasks, such as UTKFace (Zhang et al., 2017), Biwi Kinect (Fanelli et al., 2013), and California Housing (Nugent, 2017). The results showed that our SSA retains the important feature subspace during TTA and outperforms existing TTA baselines that were originally designed for classification by aligning the feature subspace.
Table 1: Number of valid (having non-zero variance) feature dimensions and feature subspace dimensions (i.e., the rank of the feature covariance matrix), and $R^{2}$ scores on the test datasets. Although the original feature space has 2048 dimensions in the experiment except for the California Housing dataset, the features of the regression models are distributed within the subspaces that have less than a hundred dimensions. Our method improves $R^{2}$ scores by the feature subspace in most cases, whereas the naive feature alignment sometimes diverged (displayed as ‘-’) because of too few valid dimensions. See Section 4.3.1 for more details.
@@
Dataset #Valid dims. #Subspace dims. Source Naive feature alignment SSA (ours)
@@
Biwi Kinect 713 34.5 $0.706$ $0.753$ $\mathbf{0.778}$
@@
The average over a model trained on each gender/target is reported.
@@ 2 Problem Setting
We consider a setting with a neural network regression model $f_{\theta}:\mathcal{X}\to\mathbb{R}$ pre-trained on a labeled source dataset $\mathcal{S}=\{(\mathbf{x}_{i}^{\text{s}},y_{i}^{\text{s}})\in\mathcal{X}\times\mathbb{R}\}_{i=1}^{N_{\text{s}}}$, where $\mathbf{x}_{i}^{\text{s}}$ and $y_{i}^{\text{s}}$ are an input and its label, and $\mathcal{X}$ is the input space. Our goal is to adapt $f_{\theta}$ to the target domain by using an unlabeled target dataset $\mathcal{T}=\{\mathbf{x}_{i}^{\text{t}}\in\mathcal{X}\}_{i=1}^{N_{\text{t}}}$ without accessing $\mathcal{S}$. Note that the target labels $y_{i}^{\text{t}}\in\mathbb{R}$ are not available. In the source dataset $\mathcal{S}$, the data $\{(\mathbf{x}_{i}^{\text{s}},y_{i}^{\text{s}})\}$ are sampled from the source distribution $p_{\text{s}}$ over $\mathcal{X}\times\mathbb{R}$. In the target dataset $\mathcal{T}$, we assume covariate shift (Shimodaira, 2000), which is a distribution shift that often occurs in the real world. In other words, the target data $\mathbf{x}_{i}^{\text{t}}$ are sampled from the target distribution $p_{\text{t}}$ over $\mathcal{X}$ that is different from $p_{\text{s}}$, but the predictive distribution is the same, i.e., $p_{\text{s}}(\mathbf{x})\neq p_{\text{t}}(\mathbf{x})$ and $p_{\text{s}}(y|\mathbf{x})=p_{\text{t}}(y|\mathbf{x})$.
@@ 3 Test-time Adaptation for Regression
Figure 1: Overview of significant-subspace alignment (SSA). A more detailed procedure is listed in Algorithm 1.
In this section, we describe the basic idea behind Significant-subspace Alignment (SSA) in Section 3.1 and describe it in detail in Section 3.2. Our method can be applied regardless of the form of input data since SSA does not rely on input-specific method, such as image data augmentations or self-supervised tasks.
@@ 3.1 Basic Idea: Feature Alignment
$\bm{\mu}^{\text{s}}=\frac{1}{N_{\text{s}}}\sum_{i=1}^{N_{\text{s}}}\mathbf{z}_{i}^{\text{s}},\quad\bm{\sigma}^{\text{s~{}2}}=\frac{1}{N_{\text{s}}}\sum_{i=1}^{N_{\text{s}}}(\mathbf{z}_{i}^{\text{s}}-\bm{\mu}^{\text{s}})\odot(\mathbf{z}_{i}^{\text{s}}-\bm{\mu}^{\text{s}}),$ (1)
@@
Then, we move to the target domain, where we cannot access the source dataset $\mathcal{S}$. Given a target mini-batch =$\mathcal{B}=\{\mathbf{x}_{i}^{\text{t}}\}_{i=1}^{B}$ sampled from $\mathcal{T}$, we compute the mini-batch mean and variance $\hat{\bm{\mu}}^{\text{t}}$ and $\hat{\bm{\sigma}}^{\text{t~{}2}}$ analogously to Equation 1.
For feature alignment, we seek to make the target statistics similar to the source ones. For this purpose, we use the KL divergence as Nguyen et al. (2022) proved that it is included in an upper bound of the target error in unsupervised domain adaptation. Concretely, we minimize the KL divergence between two diagonal Gaussian distributions $\mathcal{N}(\bm{\mu}^{\text{s}},{\bm{\sigma}^{\text{s~{}2}}})$ and $\mathcal{N}(\hat{\bm{\mu}}^{\text{t}},\hat{\bm{\sigma}}^{\text{t~{}2}})$:
[…]
where the subscripts $d$ represent the $d$-th elements of the mean and variance vectors. Here, we used both directions of the KL divergence because it empirically had good results as recommended by Nguyen et al. (2022). The KL divergence between two univariate Gaussians can be written in a closed form (Duchi, 2007)
$D_{\text{KL}}\left(\mathcal{N}(\mu_{1},\sigma_{1}^{2})\|\mathcal{N}(\mu_{2},\sigma_{2}^{2})\right)=\left[\log(\sigma_{2}^{2}/\sigma_{1}^{2})+\{(\mu_{1}-\mu_{2})^{2}+\sigma_{1}^{2}\}/\sigma_{2}^{2}-1\right]/2.$ (3)
However, in regression models, the features tend to be less diverse than in classification (Zhang et al., 2023). In addition to the insight from Zhang et al. (2023), we observed that the features of a regression model trained on $\mathcal{S}$ are distributed in only a small subspace of the feature space and many dimensions of the feature space had zero variances. Table 1 shows the numbers of valid (having non-zero variance) feature dimensions and feature subspace dimensions (see Section 4.3.1 for more details). This property makes the naive feature alignment described above unstable since the KL divergence in Equation 3 includes the variance in the denominator. Also, this naive feature alignment is ineffective because many of the feature dimensions have a small effect on the subspace.
@@ 3.2 Significant-subspace Alignment
In this section, we describe our method, Significant-subspace Alignment (SSA), to tackle the aforementioned problem of naive feature alignment. Figure 1 shows an overview of SSA. As described in Section 3.1, the features of a regression model tend to be distributed in a small subspace of the feature space. Thus, we introduce subspace detection to detect a subspace that is representative and significant to the output and then perform feature alignment in the subspace. Subspace detection is similar to principal component analysis (PCA). Further, in our regression model, a feature $\mathbf{z}=g_{\phi}(\mathbf{x})$ is projected onto a one-dimensional line determined by $\mathbf{w}$ of $h_{\psi}$ in order to output a scalar value $\mathbf{w}^{\top}\mathbf{z}+b$. Thus, a subspace basis $\mathbf{v}^{\text{s}}_{d}$ whose direction is not orthogonal to $\mathbf{w}$ needs precise feature alignment. We use dimension weighting to prioritize such dimensions.
Subspace detection. After the training on the source dataset $\mathcal{S}$, we detect the subspace in which the source features are distributed. Instead of computing the variance of each dimension as Equation 1, we compute the covariance matrix:
@@
where the source mean vector $\bm{\mu}^{\text{s}}$ is the same as Equation 1.
@@
where $\mathbf{V}^{\text{s}}=[\mathbf{v}_{1}^{\text{s}},\ldots,\mathbf{v}_{K}^{\text{s}}]^{\top}\in\mathbb{R}^{K\times D}$. With $\tilde{\mathbf{z}}^{\text{t}}_{i}\in\mathbb{R}^{K}$, we compute the projected target mean and variance over the mini-batch analogously to Equation 1; this is denoted by ${\tilde{\bm{\mu}}^{\text{t}}}$ and $\tilde{\bm{\sigma}}^{\text{t~{}2}}$. On the other hand, the projected source mean and variance are $\mathbf{0}$ and the eigenvalues $\bm{\lambda}^{\text{s}}=[\lambda_{1}^{\text{s}},\ldots,\lambda_{K}^{\text{s}}]$ since $\bm{\Sigma^{\text{s}}}\mathbf{v}_{k}^{\text{s}}=\lambda^{\text{s}}_{k}\mathbf{v}_{k}^{\text{s}}$. Thus, the KL divergence in the detected subspace is computed between two $K$-dimensional diagonal Gaussians $\mathcal{N}(\mathbf{0},\bm{\lambda}^{\text{s}})$ and $\mathcal{N}({\tilde{\bm{\mu}}^{\text{t}}},\tilde{\bm{\sigma}}^{\text{t~{}2}})$. Using subspace detection and dimension weighting, the loss of SSA is:
@@
[…]
During TTA, we optimize the feature extractor $g_{\phi}$ to minimize $\mathcal{L}_{\text{TTA}}$, i.e., we seek $\phi^{*}=\min_{\phi}\mathcal{L}_{\text{TTA}}(\phi)$. We update only the affine parameters $\gamma$ and $\beta$ of the normalization layers such as batch normalization (Ioffe & Szegedy, 2015) or layer normalization (Ba et al., 2016) inspired by Tent (Wang et al., 2021). This strategy is effective not only to retain the source knowledge but also to enable flexible adaptation (Frankle et al., 2021; Burkholz, 2024). The procedure of SSA is listed in Algorithm 1 of the Appendix.
Is diagonal Gaussian distribution appropriate? For computing KL divergence of $\mathcal{L}_{\text{TTA}}$, we assume the source and target feature distributions as diagonal Gaussian. This is reasonable because features are likely to follow a Gaussian distribution when projected onto the feature subspace detected by subspace detection as the number of original feature dimensions increases by the central limit theorem, as described in Figure 3 and Section 4.3.4. Moreover, since subspace detection uses the PCA, the features projected onto the subspace are decorrelated. Thus, assuming that each dimension is independent, i.e., diagonal, is also reasonable.